OpenStudy (anonymous):

(x-5)(x+3)=33 ; find the real roots

OpenStudy (cwrw238):

(x-5)(x+3)=33 x^2 - 2x - 15 = 33 x^2 - 2x - 48 = 0 can u continue from here?

OpenStudy (anonymous):

Thanks! Just realized my mistake. added 33 instead of subtracting

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

Ah another question. Havent done this stuff in 4 years.

OpenStudy (anonymous):

(x+2)^2(x-1)+(x+2)(x-1)^2

OpenStudy (whpalmer4):

\[(x+2)^2(x-1) + (x+2)(x-1)^2\]multiply it out, step by step: \[(x+2)^2 = (x+2)(x+2) = x^2+2x+2x+4 = x^2+4x+4\] \[(x^2+2x+4)(x-1) = x^3-x^2+2x^2-2x+4x-4 = x^3+x^2+2x-4\] \[x^3+x^2+2x-4 + (x+2)(x-1)^2\]you do the rest, then simplify

OpenStudy (anonymous):

I've gotten down to 2x^3+3x^2-11x-2 Any help factoring that please?

OpenStudy (whpalmer4):

@cwrw238 pretty sure that those 2s outside parentheses are supposed to be ^2

OpenStudy (cwrw238):

yea - u r right

OpenStudy (cwrw238):

i'm embarrassed!!

OpenStudy (whpalmer4):

life would be easier if openstudiers couldn't copy and paste the questions, losing all of those valuable positional clues :-)

OpenStudy (cwrw238):

yea

OpenStudy (anonymous):

Guys , you didnt need to expand!

OpenStudy (whpalmer4):

we didn't need to do anything, really :-)

OpenStudy (anonymous):

just do something similar like a^2v +ab^2 = ab(a+b) then solve!

OpenStudy (anonymous):

gah im hitting myself >:(